OpenStudy (anonymous):

name the supplements for each of the following angles , 82, 10, 178, 90, 14, 45, 87, 52, or teach me how to do it myselff :]

OpenStudy (chaise):

Supplementary angles are on straight lines. They add to give 180 degrees. So to find the other angle. You take away the angle from 180 degrees. For example: 180 - 82 = 98 180 - 10 = 170. You can do the rest, let me know if you have trouble.

OpenStudy (anonymous):

so say i supplement 180 -45 =140?

OpenStudy (chaise):

180-45 is not 140. 135

OpenStudy (anonymous):

uqhhh then idk :/ math is so hard for me ;o

OpenStudy (anonymous):

180-82=102